Mdill111:

a total.of 324 tickets were sold. the number of.student.tickets was two times the number of adult tickets. .how many adult tickets were sold

jhonyy9:

write an equation using given details

Vixen:

damn idk

Vixen:

@k1ngofpadlet help?

K1NGofPadlet:

Simple algebra bro. So the number of adult tickets is x. According to the given information, the number of student tickets is twice the number of adult tickets, so we can verbalize the number of student tickets is 2x. To find the number of adult tickets, we can integrate the number of adult tickets (x) and the number of student tickets (2x) together, which gives us a total of 3x. Since we ken that the total number of tickets sold is 324, we can establish an equation: 3x = 324. To solve for x, we divide both sides of the equation by 3: x = 324/3. Calculating that gives us x = 108. Consequently, the number of adult tickets sold is 108.

KyledaGreat:

Let's solve this step-by-step: * A total of 324 tickets were sold * The number of student tickets was 2 times the number of adult tickets We want to find the number of adult tickets. Let's call the number of adult tickets x. Then the number of student tickets will be 2x. Since the total number of tickets is 324, we can write an equation: x + 2x = 324 3x = 324 Solving for x: x = 324/3 = 108 Therefore, the number of adult tickets sold was 108. In summary: * Total tickets sold = 324 * Student tickets = 2 * Adult tickets * If adult tickets = x * Then student tickets = 2x * x + 2x = 324 * 3x = 324 * x = 324/3 = 108 So, the number of adult tickets sold was 108.

K1NGofPadlet:

@kyledagreat wrote:
Let's solve this step-by-step: * A total of 324 tickets were sold * The number of student tickets was 2 times the number of adult tickets We want to find the number of adult tickets. Let's call the number of adult tickets x. Then the number of student tickets will be 2x. Since the total number of tickets is 324, we can write an equation: x + 2x = 324 3x = 324 Solving for x: x = 324/3 = 108 Therefore, the number of adult tickets sold was 108. In summary: * Total tickets sold = 324 * Student tickets = 2 * Adult tickets * If adult tickets = x * Then student tickets = 2x * x + 2x = 324 * 3x = 324 * x = 324/3 = 108 So, the number of adult tickets sold was 108.
That's great bro, but I already answered it. I don't need the assistance of AI to say the same thing that I just did. 🤷‍♂️🤦‍♂️